# Decompose-and-Compose: A Compositional Approach to Mitigating Spurious Correlation

### v1

@@ 5            ${}\mathcal{M}\leftarrow\{\}$
7                   Pick $(x^{\prime},y^{\prime})\in\mathcal{B^{\prime}}$ s.t. $y\neq y^{\prime}$ ${m}\leftarrow$ AdaptiveMasking$(f,x,y,l)$ ${m}^{\prime}\leftarrow$ AdaptiveMasking$(f,x^{\prime},y^{\prime},l)$ if  casualflag=False then
@@ 14                   $\mathcal{M}\leftarrow\mathcal{M}\cup\{(\hat{x}_{comb},y)\}$
18             $L_{total}\leftarrow L_{CE}+\alpha L_{comb}$ $\theta\leftarrow$ UpdateWeights($L_{total}$
@@ 5.1 Compared Methods
DFR [7]: They argued that even in the presence of spurious correlations, neural network classifiers still learn the core features. With this motivation, they showed that simple retraining of the last layer with group-labeled data can be sufficient to reduce the spurious correlations. We evaluated their method on our ERM trained with a ResNet50 backbone.
MaskTune [3]: Their hypothesis is the opposite of DFR. They argued that with the presence of spurious correlations, ERM models fail to learn the core features. Therefore, they fine-tune the model for one epoch using a masked version of the training data in order to force the model to focus on the core features.
CNC [40]: CnC is a contrastive learning method designed to align representations of samples within the same class that have different spurious attributes, while also distinguishing between samples of dissimilar classes that share similar spurious features.
Group DRO [22]: Group DRO aims to minimize the worst-case loss across groups with strong regularization. Their approach achieves high worst-group accuracy, but requires expensive group annotations for each training point.
JTT [11]: JTT first uses ERM to detect misclassified samples. These samples are subsequently upweighted, and the model is retrained. Their method does not require group labels.
LISA [37]: LISA is a data-augmenting technique that aims to learn invariant predictors. They intervene in samples with either the same labels but different domains or the same domains but different labels.
Table 1: A comparison of mean and worst group accuracy of several methods, including ours, on four datasets. The Group Info column shows whether each method uses group labels of train/validation data, with ✓✓indicating that group info is used in both the training and validation phase. The mean and std are reported over 3 runs on different seeds. The bold and underlined numbers indicate the best results among all methods, and methods not requiring group annotations, respectively.
@@
Base (ERM) ✗/✗ $70.8_{\pm 0.5}$ $91.6_{\pm 0.1}$ $41.7$ $96.0$ $61.3_{\pm 3.4}$ $73.9_{\pm 1.5}$ $72.8_{\pm 1.6}$ $88.5_{\pm 0.3}$
DaC-C ✗/✓ $\underline{\boldsymbol{92.6_{\pm 0.2}}}$ $94.9_{\pm 0.2}$ $76.11_{\pm 0}$ $91.35_{\pm 0.2}$ $76.0_{\pm 0.8}$ $80.0_{\pm 1.4}$ $89.0_{\pm 0.7}$ $92.2_{\pm 0.2}$
DaC ✗/✓ $92.3_{\pm 0.4}$ $\underline{\boldsymbol{95.3_{\pm 0.4}}}$ $81.9_{\pm 0.7}$ $91.4_{\pm 1.1}$ $\underline{\boldsymbol{78.3_{\pm 1.6}}}$ $\underline{\boldsymbol{79.3_{\pm 0.1}}}$ $\underline{89.2_{\pm 0.1}}$ $\underline{92.2_{\pm 0.3}}$
5.2 Datasets
In this study, We compare methods on four datasets with distribution shifts. The first three datasets are related to correlation shift and the last one includes diversity shift between the train and test sets according to the categorization introduced in [38].
Waterbirds This dataset is created by combining bird photos from the Caltech-UCSD Birds-200-2011 [29] dataset with image backgrounds from the Places dataset [42]. The birds are labeled as either waterbirds or landbirds and are placed on either water or land backgrounds. Waterbirds are more frequently shown on water backgrounds, while landbirds are more often shown on land [22].
CelebA CelebA celebrity face dataset in the presence of spurious correlations was proposed by [22]. In this dataset the binary label is assigned to the hair color and the gender is the attribute with spurious correlation with the label [13].
@@ 5.3 Setup
In addition to the main method (DaC), we test another version of our method, named DaC-C, which uses all the correctly classified samples for making combined data, and removes the hyperparameter $q$. Thus, it uses correctly classification as a way for selecting low loss samples.
For all datasets, we have trained the model in two settings: one by assuming that the model generally attends to the causal parts, and the other by the assumption that the model trained by ERM attends to the non-causal parts. For all datasets except the Dominoes, the former has better worst group accuracy on the validation set. The details for training the base ERM model and training the last layer of the models with DaC are in the Appendix.
@@ 5.4 Results
The results of our experiments along with reported results for DFR [7], Masktune [3], LISA [37], Group DRO [22], and JTT [11] on four benchmarks are illustrated in Table Tab. 1. Both the worst group accuccy and the average group accuracy as the most commonly used metrics to evaluate robustness again spurious correlation have been reported. Similar to [7], the Group Info column shows whether the label of the group (majority/minority) to which datapoints belong is available for training or validation data. Among the methods in Tab. 1, only DFR requires group info of validation data during the training phase (and not just for the model selection phase), which is shown by ✓✓.
The results of the methods annotated with *** are reproduced by our own experiments. As for the other methods, the results on the Waterbirds and CelebA datasets are from their original paper. The results for Metashift are reported by  [33]. Three methods among the compared ones, i.e. DFR, Group DRO, and LISA, need the group label during the training phase as mentioned in Tab. 1. According to these results, our method outperforms other methods that don’t require group label during training with a large margin in both mean and worst group accuracy metrics on Waterbirds, Dominoes, and Metahift datasets. Moreover, although the proposed method does not need the group label of the training data, it outperforms Group DRO and LISA on Waterbirds and Metashift datasets and is on par with DFR on these datasets. It is worth mentioning that the CelebA dataset does not match with the type of spurious correlation for which our method has been designed as mentioned in Sec. 2. More precisely, in addition to the face (including gender features) that can be considered as a non-causal part for classifying the hair color, there is also a spurious attribute for the causal part (i.e. hair) in this problem. There is a spurious correlation between the volume of the hair and the label, as will be discussed in the Appendix.
@@ 5.5.1 Effect of Combining Images
Combining the images proves to be extremely effective in enhancing the worst test group accuracy, which is evident in Fig. 5. It can be seen that the value of $\alpha\geq 1$ highly reduces the reliance on spurious patterns and also the range of the proper value of $\alpha$ is similar in different datasets.
@@ 5.5.2 Effect of the Proportion of the Selected Data
As the number of the selected samples for combining increases, more datapoint on which the base model has a higher loss will be used for combining. The model is more prone to attending to irrelevant parts in these samples, which results in, the combined images being wrongly labeled. As shown in Tab. 1, the model has the best performance according to worst group accuracy when we choose the proportion of the selected samples by hyper-parameter tuning. Selecting all the correctly-classified samples omit the need for tuning hyperparametr $q$ but it may degrade the accuracy of the method when the ERM overfits on the training samples since in this case all samples may be used for combination.
@@ 7.3 Attention-based Masking for Out-of-Distribution Generalization
8 More Training Details
8.1 ERM
@@
8.2 DaC
For all datasets, Adam optimizer with a learning rate of $0.5\times 10^{-2}$, and step learning rate scheduler with step size $5$ and gamma $0.5$ were used. The batch size was 64 for all datasets. To encourage the diversity of training data during retraining the last layer, in cases when the selected samples with low loss in each batch were only from one class, we randomly combined the selected images with others from the same class. No regularization terms were used for retraining the last layer of the model. More details regarding the number of epochs, and optimal values for $\alpha$ and $q$ are in Tab. 2. Batch size, $\alpha$ and $q$ were selected from $\{32,64\}$, $\{1,...,10\}$, and $\{0.2,0.4,0.5,0.6,0.8,1\}$ respectively, and the criteria for hyperparameter selection was the worst group accuracy on the validation set.
@@ 8.2 DaC
9 Original Masks or Inverted Ones?
As mentioned in Sec. 4.2, we train the model in two settings corresponding to the ERM casual attention and ERM non-causal attention assumptions. For the former setting, i.e. ERM casual attention, we keep the parts obtained by adaptive masking as the causal parts while for the later one, i.e. ERM non-causal attention, the parts remained by adaptive masking are considered as non-causal and thus we invert the masks in order to obtain the causal regions for DaC. Based on the worst group accuracy of the model trained by DaC on the validation set in these two settings, it can be determined whether the parts to which the model generally pays more attention are causal or non-causal. The results for both cases are in Tab. 3. According to the results, unlike the Dominoes, on Waterbirds and Metashift the model attends more to the causal components. Regarding the CelebA dataset, it seems that the attention of the model does not grasp the entire hair parts in the image, hence, the inverted mask still contains a proportion of the hair. This was also reflected in the results in Tab. 1, in which, unlike other datasets, our model has a lower performance on CelebA. For more details on the CelebA dataset, refer to Sec. 13.
@@ 12 Combined Images
13 Details on the CelebA Dataset
As mentioned in Sec. 5.4, in addition to the spurious correlation between gender (which can be inferred from the facial features) and hair colour, there are some hair attributes contributing to hair volume such as hair wave and baldness, which are correlated with the hair colour. The number of people with each hair colour and specific attributes is extracted from the CelebA metadata and shown in Tabs. 4 and 5. According to the statistics, while about $0.05\%$ of blond people wear hats or are bald, more than $8$ per cent of people who are not blond wear hats or are bald. Similarly, the percentage of blond people with wavy hair is more than $1.5$ times greater than the ones that are not blond. Additionally, our eye observations from the dataset indicate that there is a correlation between the length of hair and its colour, as short hair is more co-occurred with non-blond hair. It is worth mentioning that since the attribute of hair length was not available in the CelebA metadata, we assessed this claim by eye observation. A few examples of randomly selected samples from each hair colour are shown in Fig. 14.
Table 4: Number of people with wavy hair with each hair colour.
Blond = -1 Blond = 1
Wavy = -1 121761 16094
Wavy = 1 50855 13889
Blond = -1 Blond = 1
Bald = -1 $\land$ Hat = -1 158440 29817
Bald = 1 $\lor$ Hat = 1 14176 166

### v3

@@ 6            ${}\mathcal{M}\leftarrow\{\}$
9                   Pick $(x^{\prime},y^{\prime})\in\mathcal{B^{\prime}}$ s.t. $y\neq y^{\prime}$
10                  ${m}\leftarrow$ AdaptiveMasking$(f,x,y,l)$
11                  ${m}^{\prime}\leftarrow$ AdaptiveMasking$(f,x^{\prime},y^{\prime},l)$
12                  if  casualflag=False then
@@ 19                   $\mathcal{M}\leftarrow\mathcal{M}\cup\{(\hat{x}_{comb},y)\}$
23             $L_{total}\leftarrow L_{CE}+\alpha L_{comb}$
24            $w\leftarrow$ UpdateWeights($L_{total}$
@@ 5.1 Compared Methods
In this paper, we consider 6 baselines besides ERM. DFR [8] argues that even in the presence of spurious correlations, neural network classifiers still learn the core features. Following this, they show that simple retraining of the last layer with group-balanced data can be sufficient to make the model robust to spurious correlation. We evaluated their method on our ResNet50 [5] backbone trained with ERM. Group DRO [24] aims to minimize the worst-case loss across groups with strong regularization. LISA [39] is a data-augmenting technique that aims to learn invariant predictors By intervening on samples with either the same labels but different domains or the same domains but different labels. MaskTune [3] argues that in the presence of spurious correlations, ERM models attend more to the spurious parts of images. Therefore, they fine-tune an ERM model for one epoch using a masked version of the training data to force the model to focus on the core parts. CNC [42] is a contrastive learning method designed to align representations of samples within the same class that have different spurious attributes, while also distinguishing between samples of dissimilar classes that share similar spurious features. JTT [13] first uses a model trained with ERM to detect misclassified samples. These samples are subsequently upweighted when training a new model on the dataset.
Table 1: A comparison of mean and worst group accuracy of several methods, including ours, on four datasets. The Group Info column shows whether each method uses group labels of train/validation data, with ✓✓indicating that group info is used in both the training and validation phases. The mean and std are reported over 3 runs on different seeds. The bold and underlined numbers indicate the best results among all methods, and methods not requiring group annotation, respectively.
@@
Base (ERM) ✗/✗ $70.8_{\pm 0.5}$ $91.6_{\pm 0.1}$ $41.7$ $\underline{\boldsymbol{96.0}}$ $61.3_{\pm 3.4}$ $73.9_{\pm 1.5}$ $72.8_{\pm 1.6}$ $88.5_{\pm 0.3}$
DaC-C ✗/✓ $\underline{\boldsymbol{92.6_{\pm 0.2}}}$ $94.9_{\pm 0.2}$ $76.11_{\pm 0}$ $91.35_{\pm 0.2}$ $76.0_{\pm 0.8}$ $\underline{\boldsymbol{80.0_{\pm 1.4}}}$ $89.0_{\pm 0.7}$ $92.2_{\pm 0.2}$
DaC ✗/✓ $92.3_{\pm 0.4}$ $\underline{\boldsymbol{95.3_{\pm 0.4}}}$ $81.9_{\pm 0.7}$ $91.4_{\pm 1.1}$ $\underline{\boldsymbol{78.3_{\pm 1.6}}}$ $79.3_{\pm 0.1}$ $\underline{89.2_{\pm 0.1}}$ $\underline{92.2_{\pm 0.3}}$
@@ 5.2 Setup
The experiments are done on four datasets: Waterbirds [24], CelebA [15], Metashift [12], and Dominoes [18]. The details for these datasets are in Appendix 8.
@@
In addition to the main method (DaC), we test another version of our method, named DaC-C, which uses all the correctly classified samples for making combined data, and removes the hyperparameter $q$. Thus, it uses correct classification as a way for selecting low-loss samples.
For all datasets, we have trained the model in two settings: one by assuming that the model generally attends to the causal parts, and the other by the assumption that the model trained by ERM attends more to the non-causal parts. For all datasets except the Dominoes, the former has better worst group accuracy on the validation set.
The details for training the base ERM model and training the last layer of the model with DaC are in Appendix 8.
@@ 5.3 Results
The results of our experiments along with reported results for DFR [8], Masktune [3], LISA [39], Group DRO [24], and JTT [13] on four benchmarks are illustrated in Table Tab. 1. Both the worst group accuracy and the average group accuracy as the most commonly used metrics to evaluate robustness again spurious correlation have been reported. Similar to [8], the Group Info column shows whether the label of the group (majority/minority) to which datapoints belong is available for training or validation data. Among the methods in Tab. 1, only DFR requires group info of validation data during the training phase (and not just for model selection), which is shown by ✓✓.
The results of the methods annotated with * are reproduced by our own experiments. As for the other methods, the results on the Waterbirds and CelebA datasets are from their original paper. The results for Metashift are reported by  [35]. Three methods among the baselines, i.e. DFR, Group DRO, and LISA need the group label during the training phase as mentioned in Tab. 1. According to these results, our method outperforms other methods that don’t require group labels during training with a large margin in both mean and worst group accuracy metrics on Waterbirds, Dominoes, and Metahift datasets. Moreover, although the proposed method does not need the group label of the training data, it outperforms Group DRO and LISA on Waterbirds and Metashift datasets and is on par with DFR on these datasets. It is worth mentioning that the CelebA dataset does not match the type of spurious correlation for which our method has been designed as mentioned in Sec. 2. More precisely, in addition to the face (including gender features) that can be considered as a non-causal part for classifying the hair colour, there is also a spurious attribute for the causal part (i.e. hair) in this problem. There is a spurious correlation between the volume of the hair and the label, as will be discussed in Appendix 8.
@@ 5.4.1 Effect of Combining Images
Combining images proves to be extremely effective in enhancing the worst test group accuracy, which is evident in Fig. 5. It can be seen that the value of $\alpha\geq 1$ highly reduces the reliance on spurious patterns and also the range of the proper value of $\alpha$ is similar in different datasets.
@@ 5.4.2 Effect of the Proportion of the Selected Data
As the number of the selected samples for combining increases, more datapoint on which the base model has a higher loss will be used for combining. The model is more prone to attending to irrelevant parts in these samples, which results in the combined images being wrongly labelled. As shown in Tab. 1, the model has the best worst group accuracy when we choose the proportion of the selected samples by hyper-parameter tuning. Selecting all the correctly-classified samples omit the need for tuning hyperparameter $q$, but it may degrade the accuracy of the method when the ERM overfits on the training samples, since in this case, all samples may be used for combination.
@@ 6 Conclusion
Acknowledgments
We would like to thank Mr. Mohammad-Mahdi Samiei for his insightful and constructive comments.
@@ 7.3 Attention-based Masking for Out-of-Distribution Generalization
8 Details on Experiments
8.1 Datasets
In this study, We compared methods on four datasets with distribution shifts. The first three datasets are related to correlation shift and the last one includes diversity shift between the train and test sets according to the categorization introduced in [40].
Waterbirds This dataset is created by combining bird photos from the Caltech-UCSD Birds-200-2011 [31] dataset with image backgrounds from the Places dataset [44]. The birds are labelled as either waterbirds or landbirds and are placed on either water or land backgrounds. Waterbirds are more frequently shown on water backgrounds, while landbirds are more often shown on land [24].
CelebA CelebA celebrity face dataset in the presence of spurious correlations was proposed by [24]. In this dataset the binary label is assigned to the hair colour and the gender is the attribute with spurious correlation with the label [15].
8.2 Details on the CelebA Dataset
As mentioned in Sec. 5.3, in addition to the spurious correlation between gender (which can be inferred from the facial features) and hair colour, some hair attributes contribute to hair volume such as hair wave and baldness, which are correlated with the hair colour. The number of people with each hair colour and specific attributes is extracted from the CelebA metadata and shown in Tabs. 2 and 3. According to the statistics, while about $0.05\%$ of blond people wear hats or are bald, more than $8$ per cent of people who are not blond wear hats or are bald. Similarly, the percentage of blond people with wavy hair is more than $1.5$ times greater than the ones that are not blond. Additionally, our eye observations from the dataset indicate that there is a correlation between the length of hair and its colour, as short hair is more co-occurred with non-blond hair. It is worth mentioning that since the attribute of hair length was not available in the CelebA metadata, we assessed this claim by eye observation. A few examples of randomly selected samples from each hair colour are shown in Fig. 14.
Table 2: Number of people with wavy hair with each hair colour.
Blond = -1 Blond = 1
Wavy = -1 121761 16094
Wavy = 1 50855 13889
Blond = -1 Blond = 1
Bald = -1 $\land$ Hat = -1 158440 29817
Bald = 1 $\lor$ Hat = 1 14176 166
8.3 ERM Training Details
@@
8.4 DaC Training Details
For all datasets, Adam optimizer with a learning rate of $0.5\times 10^{-2}$, and step learning rate scheduler with step size $5$ and gamma $0.5$ were used. The batch size was 64 for all datasets. To encourage the diversity of training data during retraining the last layer, in cases when the selected samples with low loss in each batch were only from one class, we randomly combined the selected images with others from the same class. No regularization terms were used for retraining the last layer of the model. The proportions for creating the curve of the loss with respect to the amount of masking in adaptive masking did not contain 1, since masking the whole image would trivially increase the loss on the masked image significantly. More details regarding the number of epochs, and optimal values for $\alpha$ and $q$ are in Tab. 4. Batch size, $\alpha$ and $q$ were selected from $\{32,64\}$, $\{1,...,10\}$, and $\{0.2,0.4,0.5,0.6,0.8,1\}$ respectively, and the criteria for hyperparameter selection was the worst group accuracy on the validation set.
8.5 Original Masks or Inverted Ones?
As mentioned in Sec. 4.2, we train the model in two settings corresponding to the ERM casual attention and ERM non-causal attention assumptions. For the former setting, i.e. ERM casual attention, we keep the parts obtained by adaptive masking as the causal parts while for the later one, i.e. ERM non-causal attention, the parts remained by adaptive masking are considered as non-causal and thus we invert the masks in order to obtain the causal regions for DaC. Based on the worst group accuracy of the model trained by DaC on the validation set in these two settings, it can be determined whether the parts to which the model generally pays more attention are causal or non-causal. The results for both cases are in Tab. 5. According to the results, unlike the Dominoes, on Waterbirds and Metashift the model attends more to the causal components. Regarding the CelebA dataset, it seems that the attention of the model does not grasp the entire hair parts in the image, hence, the inverted mask still contains a proportion of the hair. This was also reflected in the results in Tab. 1, in which, unlike other datasets, our model has a lower performance on CelebA. For more details on the CelebA dataset, refer to Sec. 8.2.
8.6 Details on the Kneedle Algorithm
As mentioned in Sec. 5.2, we use the Kneedle algorithm for finding the optimal amount of masking in Algorithm 1. This optimal amount is indicated by the elbow (i.e. the point with the highest curvature) of the curve of the loss with respect to the amount of masking. Since we only have access to a finite number of points from this curve, we use the Kneedle algorithm, which identifies elbows in a finite set of points from a curve.
The Kneedle method is based on the concept that knee points approximate the local maxima when the set of points is rotated about a specific line. This line is determined by the first and last points and is chosen to preserve the overall behaviour of the set. By rotating the curve about this line, knee/elbow points are identified as the points where the curve deviates most from the straight line segment connecting the set’s endpoints. This approximation effectively captures the points of maximum curvature for the discrete set of points. The algorithm works as follows:
  1. 1.
    Smoothing: it applies a smoothing spline or other smoothing methods to data.
  2. 2.
    Normalization: It normalizes smoothed data by min-max normalization to function well regardless of the magnitude of data values.
  3. 3.
    Difference Computation: It defines $D_{d}$ as the set of differences between x- and y- values. The knee is where the difference curve changes from horizontal to sharply decreasing.
  4. 4.
    Local Maxima Calculation: It identifies the local maxima of the difference curve as candidate knee points.
  5. 5.
    Threshold Calculation: For each local maximum $(x_{lmx_{i}},y_{lmx_{i}})$ in the difference curve it defines the $T_{lmx_{i}}$ which is based on the average difference between consecutive $x$ values in the difference curve and a sensitivity parameter, $S$. This parameter determines how aggressive the method is. Smaller values for $S$, detect knees quicker, and large values are more conservative.
    $T_{lmx_{i}}=y_{lmx_{i}}-S.\frac{\sum_{i=1}^{n-1}x_{i+1}-x_{i}}{n-1}$
  6. 6.
    Knee Declaration: If any difference value $(x_{i},y_{i})$, where $j>i$, drops below the threshold $y=T_{lmx_{i}}$ before the next local maximum in the difference curve is reached, the method declares that local maximum as a knee point.
    Kneedle’s run time for any given n pairs of $x-$ and $y-$ values is bounded by $\mathcal{O}(n^{2})$.
8.7 Training Time
Since the ERM model used for computing the attribution scores of the pixels is fixed, extracting the attention heatmap and adaptive masking is done as a preprocess. Hence, during training, the previously prepared and saved masks are used, similar to MaskTune [3]. Additionally, since the optimal percentage of the masked pixels in Adaptive Masking is selected among a small number of candidates, the time complexity of FindElbow is constant. The training time of several methods (excluding the ERM phase of the methods) on Waterbirds is shown in Tab. 6.
Table 6: The training time of different methods (excluding the ERM training phase) on the Waterbirds dataset on Nvidia A100 GPU
Method DFR CnC JTT MaskTune Ours
Time (min) 4 85 58 6.5 18.9